A baker thinks that there is a problem with the baked bread numbers in the bakery he is working. So, he started tracking the number of breads baked in the oven. About 7 days into the bakers monitoring, he had counted 120 baked breads, after 13 days he had counted 108. a. Write the equation that represents this situation (define your variable). b. Calculate the slope in the equation (show your solution). c. Draw a graph that will represent your slope intercepts. d. What is the correlation between the number of days and the baked breads in the bakery? e. Explain the cause and effect of the slope.


1
Expert's answer
2021-06-30T19:35:02-0400

"\tilde{x}=\frac{∆xi}{2}=\frac{20}{2}=10\\\tilde{Y}=\frac{∆Yi}{2}=\frac{228}{2}=114"

Equition of the line

"Y=\beta _0+\beta _1\alpha"

Where "\tilde{\beta _1}=\frac{\varsigma(Yi-\bar{Y})(\alpha i-\bar{\sigma})}{\varsigma(\alpha I+\bar{alpha})^2}"

"\bar{\beta 1}=\frac{-36}{18}\\\bar{\beta_0}= -2\\ and \\ \bar{\beta_0}=\bar{y}=-\bar{\beta_1}\bar{\alpha}\\\bar{\beta_0}=114+2×10\\\bar{\beta_1}=134\\Y=134-2X\\R^2=1-\frac{SSRes}{SST}\\R^2=1\\=\varOmega=\pm or \plusmn 1\\\varOmega=-1"

(a)Using regression method, the equition of the line will be "Y=134-2x"

(b) The slope of the equition will be -2.

(c)


(d)The correlation is Perfectly negative correlation of -1 from the coefficient of determination.

(e)The line equation is "Y=134-2x"

X=4days

Y=baked bread

Slope is 2

There is a negative relationship between number of days and slope which means that of we increase the number of days by 1, the baked bread decreases by 2.